Grade 9Physical Chemistry

An aqueous solution containing 4% nonvolatile solute exerts a pressure of 1.002 bar at the normal boiling point of the solvent. What is the molar mass of the solute?

rishav kumar
12 Years agoGrade 9

1 Answer

Sunil Kumar FP
12 Years ago
Vapour pressure of the solution at normal boiling point (p1) = 1.004 bar

Vapour pressure of pure water at normal boiling point

Mass of solute, (w2) = 2 g

Mass of solvent (water), (w1) = 98 g

Molar mass of solvent (water), (M1) = 18 g mol−1

According to Raoult’s law,

pi0-p/pi0=w2*m1/(w1*m2)
1.013-1.004/1.013=2*18/98*m2
= 41.35 g mol−1

Hence, the molar mass of the solute is 41.35 g mol−1.